Circular No. 6453

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

NOVA SAGITTARII 1996 P. Martini, B. S. Gaudi, and R. M. Wagner, Ohio State University,

report: “Observations were obtained in the near-infrared at the Perkins 1.8-m Telescope (+ OSIRIS) on Aug. 6.27 and 8.23 UT. On Aug. 6.27, K = 8.6 and J − K = +0.2 with internal (photometric) errors of 0.1 mag. As this object was bright enough for the central pixel in each passband to enter the nonlinear region of the array, these measurements should only be considered lower limits on the object’s brightness. A K-band spectrum (range 2080–2280 nm, resolution 1.6 nm) obtained on Aug. 8.23 revealed broad Brγ emission (FWHM = 12 nm, EW = 36 nm), as well as Fe II 2089-nm and Na I 2208-nm emission lines.”

G. Sostero, S. Korica, K. Miocić, V. Rodiger, and V. Ujcić report the following photometry (accuracy ±0.05 magnitude) of N Sgr 1996 obtained with the 0.4-m telescope of the Višnjan Astronomical Observatory, using as a primary reference star the nearby star as measured by Gilmore and Kilmartin (IAUC 6443): Aug. 9.924 UT, V = 11.78, B −V = +1.00, V −R = +0.86, V −I = +1.64; 10.877, 12.04, +0.99, +0.85, +1.65; 12.855, 12.06, +0.93, +0.82, +1.61; 13.917, 11.97, +1.00, +0.80, +1.59; 14.937, 11.97, +0.98, +0.79, +1.60.

D. W. Dunham, Applied Physics Group, Johns Hopkins University; and M. Soma, National Astronomical Observatory of Japan, report that N Sgr 1996 is currently in a pattern of occultations by the moon that will continue through 1997. The next two occultation events are predicted to occur on Aug. 24 (regions bordering the Indian Ocean) and Sept. 21 (South America). Further information is available on the World Wide Web at

http://www.sky.net/~robinson/iotandx.htm. There is no known record

of the moon’s previously having been seen to occult a nova.

COMET C/1996 E1 (NEAT) Total visual magnitude estimates (cf. IAUC 6437): July 23.45 UT,

11.6 (A. Hale, Cloudcroft, NM, 0.2-m reflector); 27.05, 11.2 (K. Sárneczky, Ráktanya, Hungary, 0.44-m reflector); Aug. 4.91, 11.3 (R. J. Bouma, Groningen, The Netherlands, 0.25-m reflector); 8.26, 11.5 (C. E. Spratt, Victoria, BC, 0.20-m reflector); 10.00, 11.1 (G. Bakos, Ráktanya, Hungary, 0.44-m reflector); 12.10, 10.7 (J. Carvajal, Teruel, Spain, 0.21-m reflector).

1996 August 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green